Stocks A and B have the following probability distributions of expected future returns:

Probability A B

0.2 (9%) (22%)

0.2 5 0

0.3 11 20

0.2 20 30

0.1 40 36

A- Calculate the expected rate of return, rB, for Stock B (rA = 10.50%.) Do not round intermediate calculations. Round your answer to two decimal places. %

B- Calculate the standard deviation of expected returns, σA, for Stock A (σB = 20.02%.) Do not round intermediate calculations. Round your answer to two decimal places. %

C- Now calculate the coefficient of variation for Stock B. Round your answer to two decimal places.

D- Is it possible that most investors might regard Stock B as being less risky than Stock A?

If Stock B is more highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.

If Stock B is more highly correlated with the market than A, then it might have the same beta as Stock A, and hence be just as risky in a portfolio sense.

If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.

If Stock B is less highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be more risky in a portfolio sense.

If Stock B is more highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be less risky in a portfolio sense.
Answer #1

A. Expected Return of Stock B =0.2*-22%+0.2*0%+0.3*20%+0.2*30%+0.1*36% =11.20%

B. Standard Deviation of Stock A =(0.2*(-9%-10.50%)^2+0.2*(5%-10.50%)^2+0.3*(11%-10.50%)^2+0.2*(20%-10.50%)+0.1*(40%-10.50%)^2)^0.5 =18.73%

C. Coefficient of Variation of B =Standard Deviation/Expected return =20.02%/11.20% =1.79

d. Option III is correct option.